23.A 100.0 mL sample of 0.10 M Ca(OH)2 is titrated with 0.10 M HBr. Determine the pH of the solution before the addition of any HBr. 24. Determine the pH of the solution after the addition of 100.0 mL HBr. 25. Determine the pH of the solution after the addition of 200.0 mL HBr. 26. Determine the pH of the solution after the addition of 300.0 mL HBr. 27. Determine the pH of the solution after the addition of 400.0 mL HBr.
Answer #1

23.A 100.0 mL sample of 0.10 M Ca(OH)2 is titrated with 0.10 M HBr.

Determine the pH of the solution before the addition of any HBr.

pOH = -log(OH)

[OH-] = 2*[Ca(OH)2] = 2*(0.1) = 0.2 M

pOH = -log(0.2) = 0.698

pH = 14-0.698

pH = 13.3

24. Determine the pH of the solution after the addition of 100.0 mL HBr.

mmol of OH- = 2*MV = 2*100*0.1 = 20 mmol of OH-

mmol of H+ = MV = 0.1*V = 0.1*100 = 10 mmol of HBr

left ion = 20-10 = 10 mmol of OH-

V = 100+100 = 200 mL

[OH-] = 10/200 = 0.05 M

pOH = -log(0.05) = 1.301

pH = 14-pOH = 14-1.301

pH = 12.69

25. Determine the pH of the solution after the addition of 200.0 mL HBr.

mmol of OH- = 2*MV = 2*100*0.1 = 20 mmol of OH-

mmol of H+ = MV = 0.1*V = 0.1*200 = 20mmol of HBr

left ion = 0

this will be NEUTRAL

pH = 7

26. Determine the pH of the solution after the addition of 300.0 mL HBr.

mmol of OH- = 2*MV = 2*100*0.1 = 20 mmol of OH-

mmol of H+ = MV = 0.1*V = 0.1*300= 30 mmol of HBr

left ion = 30-20 = 10 mmol of OH-

V = 100+300 = 400 mL

[H+] = 10/400 = 0.025 M

pH= -log(0.025 ) = 1.6

27. Determine the pH of the solution after the addition of 400.0 mL HBr.

mmol of OH- = 2*MV = 2*100*0.1 = 20 mmol of OH-

mmol of H+ = MV = 0.1*V = 0.1*400= 40 mmol of HBr

left ion = 40-20 = 20 mmol of OH-

V = 100+400= 500 mL

[H+] = 20/500 = 0.04 M

pH= -log(0.04  ) = 1.39